OpenStudy (anonymous):

simplify cube root of c^12

OpenStudy (anonymous):

c^4

OpenStudy (anonymous):

(c^12)^1/3 = c^(12/3) =c ^4

OpenStudy (anonymous):

cube root of (a+b)^3

OpenStudy (anonymous):

(a+b)..proces as done by RITA